An apple falls from a tree and hits the ground 13.4 m below. With what speed will it hit the ground? The acceleration of gravity is 9.8m/s
1
Expert's answer
2019-10-02T09:38:01-0400

When the apple is above the ground at height "H" with no speed, it has total energy equal to just potential energy: "E_1 = U_1 = m g H". Just when it hits the ground, it has some speed "v", and zero potential energy, hence the total energy is "E_2 = T_2 = \frac{m v^2}{2}" . By the law of conservation of energy, "E_1 = E_2", from where "v = \sqrt{m g H} \approx 16.21 \frac{m}{s}",